"Cuthbert Industrials, Inc. prepares three-year comparative financial statements. In year 3, Cuthbert discovered an error in the previously issued financial statements for year 1. The error affects the financial statements that were issued in years 1 and 2. How should the company report the error?"

The financial statements for years 1 and 2 should not be restated; the cumulative effect of the error on years 1 and 2 should be reflected in the carrying amounts of assets and liabilities as of the beginning of year 3.

The financial statements for years 1 and 2 should be restated; an offsetting adjustment to the cumulative effect of the error should be made to the comprehensive income in the year 3 financial statements.

The financial statements for years 1 and 2 should not be restated; financial statements for year 3 should disclose the fact that the error was made in prior years.

The financial statements for years 1 and 2 should be restated; the cumulative effect of the error on years 1 and 2 should be reflected in the carrying amounts of assets and liabilities as of the beginning of year 3.

"When a company changes an accounting principle, it should report the change by reporting the cumulative effect of the change in the current year's income statement."

TRUE

FALSE

"Carla Vista Co. purchased machinery that cost $3350000 on January 4, Year 16. The entire cost was recorded as an expense. The machinery has a 9-year life and a $235000 residual value. The error was discovered on December 20, Year 18. Ignore income tax considerations. Before the correction was made, and before the books were closed on December 31, Year 18, retained earnings was understated by"

"$2,272,778.00 "

"$2,605,556.00 "

"$2,657,778.00 "

"$3,350,000.00 "
Answer #1

1.) The financial statements for Years 1 and 2 should be restated; the cumulative effect of the error on Years 1 and 2 should be reflected in the carrying amounts of assets and liabilities as of the beginning of Year 3

2.) False, it should not report effect on current year's income statement but it should report effect on previous year's income statement.

3.) $2,657,778 is correct.

Calculation :-

Retained earnings actual reduced in year 16 = $3,350,000

Retained earnings should be reduced if correct depreciation is deducted in 2 years = [(3,350,000 - 235,000)/9] x 2 = $692,222

Before corrections and before book closure in year 18, retained earnings is net understated by = 3,350,000 - 692,222 = $2,657,778